OpenStudy (anonymous):

LOGARITHMS: log42-log6/log4

OpenStudy (anonymous):

is it \[\frac{\log(42)-\log(6)}{\log(4)}\]?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

then you get \[\frac{\log(\frac{42}{6})}{\log(4)}\] as a first step

OpenStudy (anonymous):

which is \[\frac{\log(7)}{\log(4)}\] but if you want to compute this number you need a calculator

OpenStudy (anonymous):

Can you write this as \[\log _{4}7\] ?